How many calories of heat are produced, if 1 gram of a substance is completely converted into heat energy?

Grade:12

1 Answers

Subhash Kumar Mehta
5 years ago
According to Einstein theory we have E=mc^2 and m=1g=0.001kg so E=0.001*3*10^8 =3*10^5 Ans. »Thanks»Regards»Subhash Kumar»St. Xavier`s College, Ranchi